Find the general solution.

14. z(t) = (+5 -1)=44
Answer #1

14. Given that x(t) = x(). 4) = (5 %) 2013 264) () net x(t) = Then dhe de 2 xi -5 dacz dr 22 .: dze, = लेन X1+ 2x2 D24 - x -11) is The general solution of *4 = C, Cos (3+) + C, Sin (34) from we have 2.X2 = (0-1) 21 =D24-24 doy df -X :, 2x2 =-34 Sin(

I think it will help you . If you have any doubts feel free to ask in the comments section .